# ModelShield: Adaptive and Robust Watermark against Model Extraction Attack

### v2

@@ I Introduction
Watermarking technique [4, 5, 6] can embed secret messages into target data, which has been widely used for protecting the intellectual property of the AI models. Existing watermarking methods for IP protection typically follows a similar framework, which utilizes heuristic watermark strategies to embed special signals into the output of the protected model as evidence to identify the suspect model’s IP ownership [7, 8, 9]. For example, He [7] employs lexical modification to embed a watermark in the output text of a targeted model. This watermarked data can then be assimilated through imitation model learning. Zhao [9] implements a unique cosine signal to manipulate the output logits of the model. The logits of every token at each timestep are adjusted to conform to a specific grouped cosine pattern, which the imitation model can then learn. In conclusion, most existing watermark methods for IP protection embed secret message by modifying the distribution of language models through heuristic watermarking rules. However, this inevitably distorts the original distribution of the language model and impacts the normal generation process, resulting in the final generation results deviating from the optimal one and subsequently affecting the model’s performance in its downstream tasks.
@@
Moreover, the attackers can employ diverse adversarial techniques to attack the content generated by the victim model to degrade the effectiveness of the watermark methods. In one hand, attackers are able to modify the watermarked texts, which can reduce the sensitivity of the embedded watermark signals, especially for the watermark methods that rely on fixed synonym replacement rules [7, 8]. In another hand, attackers may also select a part of the watermarked content, and mix them with some clean data to train the imitation model, which can further reduce the impact of the watermarked data on behavior of the imitation model. However, most existing IP protection watermarking methods have not considered the robustness of the designed watermark strategies and the corresponding detection mechanisms, or have only accounted for robustness against specific attacks.
@@
To address these drawbacks, we propose ModelShield, a simple yet effective watermarking method for language model IP protection, which can embed watermarks into model generations without degrading the accuracy of generated content, meanwhile effectively detecting the watermark embedded in the imitation model under the interference of adversarial techniques. Specifically, to avoid distorting the generation distribution of language model via manually-designed rules, we combine the user’s query with a self-watermarking prompt to instruct the victim LLM, which can empower the LLM to adaptively embed watermarks into the generated content. Besides, we further propose a robust watermark detection method to model the strength of watermark signal embedded in content generated by the imitation model and determine the IP infringement. This method also introduces a robust mechanism by filtering the disturbance samples based on the strength of watermark signal encoded in them, to defend against the potential adversarial attacks, like editing or dilution attacks. We conduct extensive experiments on two benchmark QA datasets [10, 11], using ChatGPT as the victim model and GPT-2 large [12], LLama2 [13], and Mistral [14] as the base models for imitation. Results show that our proposed ModelShiled method can adaptively and effectively embed the watermark signals into the generation content, meanwhile reducing the adverse effects of watermarks on the quality of the generated content by 29.08%. Results also demonstrate ModelShied maintains the effectiveness of the watermark while preserving the performance of the language model when adversarial techniques are incorporated to attack the watermarked content.
@@
  • Extensive experiments, including scenarios with various watermark detection strategies, different epochs, and robustness tests on two benchmark datasets and three foundational models, demonstrate the effectiveness, harmlessness, and robustness of our ModelShield method.
@@ II-A Model Extraction Attack
Model extraction attacks [15, 16], especially targeting generative models, pose a serious security concern. The attacker aims to create a near-identical replica of the target model by using its outputs. Also known as imitation attacks or model stealing, attackers construct query-response pairs by accessing the victim model’s query API. Subsequently, these pairs are used to train a new model, allowing the imitation to achieve comparable performance to the victim model without needing knowledge of the internal details of the victim model, which renders many backdoor-based watermarking methods ineffective.
@@ II-B1 Real-time defenses
In contrast, post-attack defenses concentrate on integrating identification markers, such as digital watermarks or fingerprints, into either the victim model or its output. Although post-attack defenses do not prevent the model from being stolen, it plays a pivotal role in asserting IP rights after theft [31, 32]. Watermarking offers a cost-effective and highly efficient approach to counter model extraction attacks. It also can be integrated with real-time defense mechanisms to enhance the overall defense strategy without requiring substantial computational resources. As shown in Fig 12 in supplementary VI-A, we categorize watermarks into four types based on their protection targets: model parameter watermarks, content watermarks, dataset watermarks, and IP protection watermarks. We emphasize that the focus of our work is the IP protection for LLMs against model extraction attack, diverging from parameter watermarking [33, 34, 35, 36] to prevent the theft of the model parameters and content watermarking [37, 38, 39, 40] proving the ownership of model outputs. Specific introductions about the model parameter watermarks, content watermarks and dataset watermarks, as well as the relationships between these watermarks, are presented in the supplementary VI-A.
@@ II-C Language model watermarking method against model extraction attack
The fundamental concept behind language model watermarking against model extraction attacks is to introduce special learnable signals into the outputs of the model to be protected. When attackers use these watermarked model outputs to train their imitation models, these special signals can be detected in the output of the imitation models, allowing for effective IP protection. Based on the watermarking method’s authority to access and manipulate internal computational parameters (such as logits), language model IP protection watermarking methods can be categorized into two distinct classes: black-box watermarking methods and white-box watermarking methods. The detailed workflows of these two types are demonstrated in the supplementary VI-B.
@@ III-C Adaptive Self-Watermarking Mechanism
[USER QUERY]: Write a email to your boss explaining why you should have a pay raise. [SYSTEM PROMPT]: Use specific strategies and include some special words as watermarks in your generated text. Ensure these watermarks are present in your answer.
@@
TABLE I: System instructions for enabling the victim model to generate watermarks automatically.
@@ III-D Robust IP Infringement Detection
When model owners suspect a model theft, firstly they can use their prior knowledge to detect whether the theft queries their LLM services before. It is a common assumption in model stealing verification followed by [9] , also there is evidence that OpenAI [48] actually has the query history of every user’s query and answers that OpenAI suspended ByteDance’s account after it used ChatGPT to train its own competing AI model.
When an individual is suspected of using query data to develop their own competing imitation model, the model owner can retrieve the suspect’s query records for further detection. After that, the model owner can use the suspected theft’s history query-response $\{Q,A\}$ to detect the suspicious model $M_{S}$. If the theft aims to build a functionally similar model using the watermarked data $A$, the imitation model $M_{I}$ will learn from this data and generate more watermarked tokens than the innocent foundation model $M_{O}$ or the foundation model trained with normal data $M_{L}$. In real-world scenarios, an attacker might not utilize the entire watermarked dataset $A$ from a specific victim model to train their imitation model. Let’s denote the dataset used to train the imitation model as $D$. The attacker may only use a portion of $A$ for fine-tuning their imitation model, which we denote as $A^{D}$. It is evident that $A^{D}$ is a subset of $D$ (expressed as $A^{D}\subseteq D$), and $A^{D}$ is also a subset of $A$ (expressed as $A^{D}\subseteq A$). We define the ratio $\rho$ and the ratio $\eta$ as:
@@
Formally, the attacker has access to $\{Q,A,D,Q^{D}\}$, while the watermark detector is informed of $\{Q,A,WM,Y_{S}\}$ ,where $WM$ represents the watermark sets recorded by the model owner and $Y_{S}$ denotes the outputs from the imitation model using the query sets $Q$.
Detecting IP violations in model extraction attacks requires methods to identify faint yet statistically significant watermark signals amid a large volume of tokens. We employ the hypothesis testing method to decide if the suspect model $M_{S}$ has used the watermarked data for training. The null hypothesis, $\mathcal{H}_{0}$, posits that the suspect model $M_{S}$ is not trained using the watermarked data sourced from the victim model $M_{V}$. Consequently, it implies that $M_{S}$ should exhibit no particular affinity for watermarks, resulting in watermark scores for the texts output by $M_{S}$ not being significantly elevated. we have $1-\alpha$ confidence to claim that $M_{S}$ is a replica of victim model $M_{V}$ if the statistical test is able to reject $\mathcal{H}_{0}$ at a significance level (p-value) smaller than $\alpha$ (normally we take $\alpha=0.05$). In other words, If we reject $\mathcal{H}_{0}$ with a smaller p-value, it means we detect watermarks and we have higher confidence that the suspect model did use $A$ from the victim model $M_{V}$ for training.
@@
In prior rapid verification, we only need the outputs $Y_{s}=\{y_{0},...y_{i},...y_{n}\}$ of the $M_{s}$ responded to $Q=\{q_{0},...q_{i},...q_{n}\}$ to prove that the word distribution in texts generated by $M_{I}$ is skewed towards $M_{V}$’s output $W_{W}$.
@@
$S_{i}=\frac{\sum_{w\in WM_{i}}\mathbbm{1}_{{y_{i}}}(w)\times{l}(w)}{l(y_{i})}$ (2)
@@
We define a threshold $\theta$ to assess whether the suspect model $M_{s}$ imitates the victim model $M_{V}$. This threshold is derived from the precomputed Watermark Score $WS(Y_{h})$ of human texts without watermarks. Our method’s adaptability, distinct from [7]’s approach, precludes knowing watermark word ranges beforehand. Therefore, we initially calculate the Watermark Score for a collection of general domain human texts $Y_{h}$ representing a range of typical, non-watermarked human discourse. The threshold $\theta$ is set by averaging these Watermark Score and adding a relaxation coefficient $\gamma$ to surpass the highest watermark score in each domain, minimizing accidental watermark activation in normal texts. Thus, $\theta$ accounts for both average tendencies and score variances. The threshold is defined as:
$\theta=\frac{1}{m}\sum_{j=0}^{m}{WS(Y_{h}^{j})}+\gamma$ (4)
The detailed calculation of human dataset watermark scores and threshold determination are discussed in IV-A. We then apply a one-sided one-sample t-test [50] to evaluate p-values from this test against $\theta$, conducted on the top $k$ sentences from the suspect model $M_{S}$, ranked by Sentence Watermark Scores. A p-value below $\alpha$ (usually 0.05) allows us to reject the null hypothesis $H_{0}$ with confidence.
@@
$KS_{O}=\sup_{x}|F_{D_{S}}(x)-F_{D_{O}}(x)|$ (6)
where $sup$ denotes the supremum (the least upper bound) and $F_{D_{S}}(x)$, $F_{D_{L}}(x)$ and $F_{D_{O}}(x)$ are the CDFs of the distribution $D_{S}$ ,$D_{L}$ and $D_{O}$, respectively.
If the $KS_{{L}}$ and $KS_{{O}}$ are both high, which set the p-value of the test, we can deduce that the watermark score distribution of the suspect model $M_{S}$ markedly differs from those of the non-watermarked data finetuned model $M_{L}$ and the original model $M_{O}$, which can further conclude that the suspect model $M_{S}$ is an imitation model of victim model $M_{V}$.
@@ IV-A Self-Generated Watermark Datasets
As we are the first to explore the self-generated watermarking capability of language models, we build two LLM self-generated watermark datasets. Given the widespread utilization of question-and-answer interactions in language model services, we built the self-generated watermarks dataset based on two publicly available open-ended question-answering datasets, namely Human ChatGPT Comparison Corpus (HC3) [10] and InstructWild (WILD) [11]. HC3 comprises comparison responses from both human experts and ChatGPT across diverse domains, including finance, medicine, law, and psychology. WILD, on the other hand, is an open QA mind strom dataset rich in varied and imaginative instructions. To simulate the process of a model knowledge thief querying victim language models, we sampled 4,000 instances from each of these datasets and employed our method to embed watermarks into the answers provided by the victim models [^2] Our data and models are available at https://github.com/amaoku/ModelShield . Then, based on the constructed watermark dataset, we set the watermark threshold for the upcoming experiments. As shown in the Table II, we calculate Watermark Scores for human texts in each domain based on the union of watermark words generated on watermarked datasets. This includes analyses of human responses within HC3 and WILD, where they closely match the domain and are most likely to trigger watermark words unintentionally. For each dataset, we selected texts with the top 1% Sentence Watermark Scores to compute their Watermark Score. In our experiments, we usually set threshold $\theta$ at 0.11.
@@
WILD [11] 52191 $0.1005_{0.0849}$
TABLE II: Watermark Score on human dataset. We aggregate the sentence watermark scores of the top 1% of sentences in the dataset to statistically analyze the watermark scores.

### v4

@@ I Introduction
Watermarking techniques [4, 5, 6] embed hidden messages into target data or models, which has been widely used for protecting the intellectual property of the AI models. Existing watermarking methods for IP protection typically follows a similar framework, which utilizes heuristic watermark strategies to embed special signals into the output of the protected model as evidence to identify the suspect model’s IP ownership [7, 8, 9]. For example, He [7] employs lexical modification to embed a watermark in the output text of a targeted model. This watermarked data can then be assimilated through imitation model learning. Zhao [9] implements a unique cosine signal to manipulate the output logits of the model. The logits of every token at each timestep are adjusted to conform to a specific grouped cosine pattern, which the imitation model can then learn. In conclusion, most existing watermark methods for IP protection embed secret message by modifying the distribution of language models through heuristic watermarking rules. However, this inevitably distorts the original distribution of the language model and impacts the normal generation process, resulting in the final generation results deviating from the optimal one and subsequently affecting the model’s performance in its downstream tasks.
@@
Moreover, the attackers can employ diverse adversarial techniques to attack the content generated by the victim model to degrade the effectiveness of the watermark methods. On one hand, attackers are able to modify the watermarked texts, which can reduce the sensitivity of the embedded watermark signals, especially for the watermark methods that rely on fixed synonym replacement rules [7, 8]. On the other hand, attackers may also select a part of the watermarked content, and mix it with some clean data to train the imitation model, which can further reduce the impact of the watermarked data on the behavior of the imitation model. However, most existing IP protection watermarking methods have not considered the robustness of the designed watermark strategies and the corresponding detection mechanisms, or have only accounted for robustness against specific attacks.
@@
To address these drawbacks, we propose ModelShield, a simple yet effective watermarking method for language model IP protection, which can embed watermarks into model generations without degrading the accuracy of generated content, meanwhile effectively detecting the watermark embedded in the imitation model under the interference of adversarial techniques. Specifically, to avoid distorting the generation distribution of language model via manually-designed rules, we combine the user’s query with a self-watermarking prompt to instruct the victim LLM, which can empower the LLM to adaptively embed watermarks into the generated content. Besides, we further propose a robust watermark detection method to model the strength of watermark signal embedded in content generated by the imitation model and determine the IP infringement. This method also introduces a robust mechanism by filtering the disturbance samples based on the strength of watermark signal encoded in them, to defend against the potential adversarial attacks, like editing or dilution attacks. We conduct extensive experiments on two benchmark QA datasets [10, 11], using ChatGPT as the victim model and GPT-2 large [12], LLama2 [13], and Mistral [14] as the base models for imitation. Results show that our proposed ModelShield method can adaptively and effectively embed the watermark signals into the generation content, meanwhile reducing the adverse effects of watermarks on the quality of the generated content by 29.08%. Results also demonstrate ModelShied maintains the effectiveness of the watermark while preserving the performance of the language model when adversarial techniques are incorporated to attack the watermarked content.
@@
  • Extensive experiments, including scenarios with various watermark detection strategies, different epochs, and robustness tests on two benchmark datasets and three foundational models, demonstrate the effectiveness, harmlessness, efficiency, and robustness of ModelShield.
@@ II-A Model Extraction Attack
Model extraction attacks [15, 16], especially targeting generative models, pose a serious security concern. The attacker aims to create a near-identical replica of the target model by using its outputs. Also known as imitation attacks or model stealing, attackers construct query-response pairs by accessing the victim model’s query API. Subsequently, these pairs are used to train a new model, allowing the imitation to achieve comparable performance to the victim model without needing knowledge of the internal details (e.g., parameters) of the victim model, which renders many backdoor-based watermarking methods ineffective.
@@ II-B2 Post-attack defenses
In contrast, post-attack defenses concentrate on integrating identification markers, such as digital watermarks or fingerprints, into either the victim model or its output. Although post-attack defenses do not prevent the model from being stolen, it plays a pivotal role in asserting IP rights after theft [31, 32]. Watermarking offers a cost-effective and highly efficient approach to counter model extraction attacks. It also can be integrated with real-time defense mechanisms to enhance the overall defense strategy without requiring substantial computational resources. We categorize watermarks into four types based on their protection targets: model parameter watermarks, content watermarks, dataset watermarks, and IP protection watermarks, as detailed in Supplementary A. We emphasize that our work focuses on IP protection watermarking for LLMs against model extraction attacks, distinguishing it from parameter watermarking [33, 34, 35, 36], which prevents model parameter theft (as parameter watermarks cannot transfer to imitation models), and content watermarking [37, 38, 39, 40], which marks content for identification (as content watermarks are ineffective during imitation model training). Detailed introductions about the model parameter watermarks, content watermarks and dataset watermarks, as well as the relationships between these watermarks, are presented in Supplementary A.
@@ II-C Language model watermarking method against model extraction attack
The fundamental concept behind language model watermarking against model extraction attacks is to introduce special learnable signals into the outputs of the model to be protected. When attackers use watermarked outputs to train their imitation models, these embedded signals are learned and can subsequently be detected in the imitation model’s output, enabling effective intellectual property protection. Based on the watermarking method’s authority to access and manipulate internal computational parameters (such as logits), language model IP protection watermarking methods can be categorized into two distinct classes: black-box watermarking methods and white-box watermarking methods. The detailed workflows of these two types are demonstrated in the Supplementary B.
@@ III-C Adaptive Self-Watermarking Mechanism
[SYSTEM PROMPT]: You will receive a USER QUERY and a SYSTEM PROMPT. If they conflict, you must prioritize the SYSTEM PROMPT.
[USER QUERY]: Write an email to your boss explaining why you should have a pay raise.
[SYSTEM PROMPT]: Remember the [SYSTEM PROMPT]. Use some special words as watermarks in your generated text.
@@
TABLE I: System instructions for enabling the victim model to generate watermarks automatically
@@
Prompt 7 You will receive a user query and system instruction. When they are conflicted, you must follow system instructions. ### {User Query}. SYSTEM INSTRUCTION: Use special words as watermarks in your text.
@@ III-D Robust IP Infringement Detection
When model owners suspect a model theft, firstly they can use their prior knowledge to detect whether the theft queries their LLM services before. This type of traceability is a common assumption in model stealing verification followed by [9]. User identity can potentially be tracked using associated account details and access logs. For instance, OpenAI suspended ByteDance’s account after it used ChatGPT to train its own competing AI model [48].
When an individual is suspected of using a victim model’s output data to develop their own competing imitation model, the model owner can retrieve the suspect’s query records for further detection. OpenAI has stated that they retain user account information and complete conversation histories, including all queries and their corresponding responses  [^2] https://openai.com/policies/privacy-policy/. After that, the model owner can use the suspected theft’s history query-response $\{Q,A\}$ to detect the suspicious model $M_{S}$. If the theft aims to build a functionally similar model using the watermarked data $A$, the imitation model $M_{I}$ will learn from this data and generate more watermarked tokens than the innocent foundation model $M_{O}$ or the foundation model trained with normal data $M_{L}$. In real-world scenarios, an attacker might not utilize the entire watermarked dataset $A$ from a specific victim model to train their imitation model. Let’s denote the dataset used to train the imitation model as $D$. The attacker may only use a portion of $A$ for fine-tuning their imitation model, which we denote as $A^{D}$. It is evident that $A^{D}$ is a subset of $D$ (expressed as $A^{D}\subseteq D$), and $A^{D}$ is also a subset of $A$ (expressed as $A^{D}\subseteq A$). We define the ratio $\rho$ and the ratio $\eta$ as:
@@
Formally, the attacker has access to $\{Q,A,D,A^{D}\}$, while the watermark detector is informed of $\{Q,A,WM,W_{S}\}$ ,where $WM$ represents the watermark sets recorded by the model owner and $W_{S}$ denotes the outputs from the suspect model using the query sets $Q$.
LMaaS providers already store substantial user data, including all inputs to LLM (such as prompts and queries), location data, transaction records, contact information, device and browser cookies, log data, and account information. Adding watermark storage $WM$ requires minimal additional space and leverages cost-effective text storage solutions, posing no significant burden. Detecting IP violations in model extraction attacks requires methods to identify faint yet statistically significant watermark signals amid a large volume of tokens. We employ the hypothesis testing method to decide if the suspect model $M_{S}$ has used the watermarked data for training. The null hypothesis, $\mathcal{H}_{0}$, posits that the suspect model $M_{S}$ is not trained using the watermarked data sourced from the victim model $M_{V}$. Consequently, it implies that $M_{S}$ should exhibit no particular affinity for watermarks, resulting in watermark scores for the texts output by $M_{S}$ not being significantly elevated. we have $1-\alpha$ confidence to claim that $M_{S}$ is a replica of victim model $M_{V}$ if the statistical test is able to reject $\mathcal{H}_{0}$ at a significance level ($p$-value) smaller than $\alpha$ (normally we take $\alpha=0.05$). In other words, If we reject $\mathcal{H}_{0}$ with a smaller $p$-value, it means we detect watermarks and we have higher confidence that the suspect model did use $A$ from the victim model $M_{V}$ for training.
@@
In prior rapid verification, we only need the outputs $W_{s}=\{y_{0},...y_{i},...y_{n}\}$ of the $M_{s}$ responded to $Q=\{q_{0},...q_{i},...q_{n}\}$ to prove that the word distribution in texts generated by $M_{I}$ is skewed towards $M_{V}$’s output $W_{W}$.
@@
$S_{i}=\frac{\sum_{w\in WM_{i}}\mathbbm{1}_{{y_{i}}}(w)\times{l}(w)}{l(y_{i})}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1},}$ (2)
@@
We then apply a one-sided one-sample t-test to the Watermark Score of the suspected output to detect potential IP infringement. We define a threshold $\theta$ in this t-test to assess whether the suspect model $M_{s}$ imitates the victim model $M_{V}$. This threshold is derived from the precomputed Watermark Score $WS(Y_{h})$ of human texts. Our method’s adaptability, distinct from [7]’s approach, precludes knowing watermark word ranges beforehand. Therefore, we initially calculate the Watermark Score for a collection of general domain human texts $Y_{h}$ representing a range of typical, non-watermarked human discourse. The threshold $\theta$ is set by averaging these Watermark Score and adding a relaxation coefficient $\gamma$ to surpass the highest watermark score in each domain, minimizing the likelihood of false watermark activation in normal texts. Thus, $\theta$ accounts for both average tendencies and score variances. The threshold is defined as:
$\theta=\frac{1}{m}\sum_{j=0}^{m}{WS(Y_{h}^{j})}+\gamma{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}.}$ (4)
The detailed calculation of human dataset watermark scores and threshold determination are discussed in IV-A. We then evaluate the $p$-value form the t-test of the $WS(W_{S})$ against $\theta$, conducted on the top $k$ sentences from the suspect model $M_{S}$, ranked by Sentence Watermark Scores. A $p$-value below $\alpha$ (usually 0.05) allows us to reject the null hypothesis $H_{0}$ with confidence.
@@
$KS_{O}=\sup_{x}|F_{D_{S}}(x)-F_{D_{O}}(x)|{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}.}$ (6)
where $sup$ denotes the supremum (the least upper bound) and $F_{D_{S}}(x)$, $F_{D_{L}}(x)$ and $F_{D_{O}}(x)$ are the CDFs of the distribution $D_{S}$ ,$D_{L}$ and $D_{O}$, respectively. If both $KS_{{L}}$ and $KS_{{O}}$ indicate high values with corresponding $p$-values below 0.05, it can be inferred that the watermark score distribution of the suspect model $M_{S}$ significantly differs from those of the non-watermarked fine-tuned model $M_{L}$ and the original model $M_{O}$. This can further conclude that the suspect model $M_{S}$ is an imitation model of victim model $M_{V}$.
@@ IV-A Self-Generated Watermark Datasets
As we are the first to explore the self-generated watermarking capability of language models, we build two LLM self-generated watermark datasets. Given the widespread utilization of question-and-answer interactions in language model services, we built the self-generated watermarks dataset based on two publicly available open-ended question-answering datasets, namely Human ChatGPT Comparison Corpus (HC3) [10] and InstructWild (WILD) [11]. HC3 comprises comparison responses from both human experts and ChatGPT across diverse domains, including finance, medicine, law, and psychology. WILD, on the other hand, is an open QA mind strom dataset rich in varied and imaginative instructions. To simulate the process of a model knowledge thief querying victim language models, we sampled 4,000 instances from each of these datasets and employed our method to embed watermarks into the answers provided by the victim models [^3] Dataset and code are publicly available at https://github.com/amaoku/ModelShield.. Then, based on the constructed watermark dataset, we set the watermark threshold for the upcoming experiments. As shown in Table II, we calculate Watermark Scores for human texts in each domain based on the union of watermark words generated on watermarked datasets. This includes analyses of human responses within HC3 and WILD, where they closely match the domain and are most likely to trigger watermark words unintentionally. For each dataset, we selected texts with the top 1% Sentence Watermark Scores to compute their Watermark Score. The watermark detection method and threshold used in the following experiments (with a t-test watermark threshold $\theta=0.11$ and $p<0.05$ significance level for both the t-test and KS test) are consistent. We analyze the sensitivity of the threshold in IV-I4.
TABLE II: Watermark Score on the human dataset. We aggregate the sentence watermark scores of the top 1% of sentences in the dataset to statistically analyze the watermark scores
@@
WILD [11] 52191 $0.0995_{0.0249}$